# Plug-and-Hide: Provable and Adjustable Diffusion Generative Steganography

### 03 Mar 2026 v4

@@ 5.3 Method Comparison
GSF[26] 25.02 91.98% 51.87% 44.17 73.47% 99.73% 51.80 62.85% 99.87%
GSN[25] 16.11 74.31% 99.80% 17.02 60.75% 99.47% 19.98 52.25% 99.54%
S2IRT[35] 44.52 50.10% 99.85% 45.74 50.47% 99.90% - - -
PA-B2G1(.08) 7.70 90.81% 51.75% 7.68 80.89% 52.45% 7.72 72.88% 50.90%
Table 3: Comparison with GAN-based and Flow-based GS methods using metrics FID, $Acc_{s}$, and, $\overline{Acc}$ on CelebA $64\times 64$ under various payloads. PA-B2G1(.08) denotes PA-B2G in mode I with $\Delta_{g}=0.08$.
Comparison with Prior DM-GIS Methods. We compare PA-B2G with message mapping methods including MN, MB, MC, and Multi-bits proposed in [8], GSD [27], and Gaussian Shading (GS) [29]. All message mapping methods are evaluated on CIFAR-10 $32\times 32$, FFHQ $64\times 64$, and LSUN-Bedroom $256\times 256$ datasets. For each dataset, we equip each mapping method with the same generative backbone and sampling scheduler. Specifically, we employ the $2^{\text{nd}}$ order Heun ODE solver with 20 sampling steps on CIFAR-10 $32\times 32$ and with 42 sampling steps on FFHQ $64\times 64$, while using the $1^{\text{st}}$ order ODE solver proposed in DDIM [21] with 18 sampling steps on LSUN-Bedroom $256\times 256$. We set the error tolerance $e=0.0185$ defined in Algorithm˜3 for CIFAR-10 $32\times 32$ and FFHQ $64\times 64$, and $e=0.007$ for LSUN-Bedroom $256\times 256$. Before message hiding, we pad all secret bits and then encrypt them to meet required payloads. All mapping methods utilize the same random seed for stego image generation, and the corresponding generated images are saved in PNG format. Note that the payloads listed in Table˜2 are effective payloads.
@@
Comparison with GAN- and Flow-based Methods. We also compare PA-B2G with two flow-based methods, GSF [26] and S2IRT [35], and a GAN-based method, GSN [25], on the CelebA $64\times 64$ dataset with payloads ranging from 3 to 9 bpp. GSF and S2IRT utilize the same pre-trained Glow model. For S2IRT, to maintain computational efficiency, we set the group number $K$ to 30 for a 3 bpp payload and to 40 for 6 bpp. At higher payloads, S2IRT requires a larger $K$, making it too time-consuming to generate 50,000 stego images; thus, we do not report its corresponding results in Table˜3. Given that PA-B2G in Mode I is superior to Mode II in extraction accuracy, we select it with $\Delta_{g}=0.08$ and $e=0.007$ for comparison. We employ the pre-trained diffusion model from DDIM [21] but replace its sampling scheduler with a Heun ODE solver, as suggested in Section˜4.4. As shown in Table˜3, PA-B2G significantly outperforms GAN- and flow-based methods, particularly in image quality and security, primarily due to the powerful generation capabilities of diffusion models and our provable, adjustable message mapping.
@@
Figure 3: Visualization results of the adjustment strategies adopted by [8] and PA-B2G on CIFAR-10, FFHQ, and LSUN-Bedroom.
@@
Figure 4: Visualization results of combining PA-B2G in mode I with Stable Diffusion on text-to-image and image-to-image tasks.
@@
Comparison with Prior Adjustment Strategies. Kim et al. [8] identify a trade-off between stego image quality and extraction accuracy, proposing MN, MB, and MC as balancing options. To evaluate our adjustment strategy (Algorithm˜3), Table˜2 reports the FID score, $\overline{Acc}$, and $Acc_{s}$ for these methods alongside PA-B2G (modes I &\& II) across varying $\Delta_{g}$. For better intuition, Figure˜3 illustrates further adjustment results, confirming that image quality, security, and accuracy can be effectively balanced by tuning $\Delta_{g}$. This supports our analysis in Section˜4.1 and reveals that Mode I provides a more natural, fine-grained adjustment than [8]. By modulating $\Delta_{g}$, PA-B2G can readily match the performance of MN, MB, and MC, as they represent special cases of our generalized framework.
5.4 Diffusion Model Watermarking
JPEG Random Crop Gaussian Noise Gaussian Blur
Payload 90 75 50 0.1 0.3 0.5 0.01 0.03 0.05 2 3 4
256 bits 100.00% 99.89% 99.46% 99.98% 99.04% 87.05% 99.85% 98.03% 94.45% 99.94% 99.25% 96.64%
512 bits 99.80% 99.27% 98.20% 99.51% 94.20% 78.25% 98.79% 94.95% 85.58% 99.26% 96.58% 91.48%
1024 bits 98.60% 97.30% 94.45% 97.48% 87.88% 71.22% 96.64% 88.86% 81.93% 96.04% 90.76% 83.43%
@@ 5.4 Diffusion Model Watermarking
Table 4: Robustness evaluation of average message extraction accuracy (%) under various image distortions (JPEG compression with different quality factors, random cropping with varying ratios, Gaussian noise with varying standard deviations, and Gaussian blurring with varying radii) across different payload sizes (256, 512, and 1024 bits)
@@
In this section, we further validate the effectiveness of PA-B2G in model I for watermarking diffusion models. We employ Dreamshaper 7, which is fine-tuned on Stable Diffusion 1.5, as the generative backbone, and utilize a first-order DDIM solver with 50 sampling steps for stego image generation. We evaluate three effective payloads—256, 512, and 1024 bits—and report their corresponding average message extraction accuracy ($\overline{Acc}$) over 1000 stego images against various lossy processing techniques in Table˜4. These techniques include JPEG compression with different quality factors, random cropping with varying ratios, Gaussian noise with varying standard deviations, and Gaussian blurring with varying radii. For PA-B2G, the error tolerance $e$ is set to 0.02, code length $l$ to 1, and $\Delta_{g}$ to 0.05.
@@
As reported in Table˜4, PA-B2G exhibits robust resistance to nearly all lossy processing when the effective payload is 256 bits. Notably, even with a cropping ratio of 0.5, PA-B2G still guarantees an extraction accuracy of 87.05%. While the extraction accuracy gradually decreases as the payload increases, PA-B2G with a 1024-bit payload still effectively resists JPEG compression ($\overline{Acc}>94\%$), which is a common operation on social media platforms. In addition to extraction accuracy, we present the visual generation results for text-to-image and image-to-image tasks in Figure˜4. As shown, the stego images produced by PA-B2G exhibit visual quality indistinguishable from normally generated images. These experimental results demonstrate that PA-B2G can be effectively extended to the watermarking of diffusion models.
$32\!\times\!32\!\times 3$ $64\!\times\!64\!\times 3$ $128\!\times\!128\!\times 3$
Method 3 bpp 6 bpp 3 bpp 6 bpp 3 bpp 6 bpp
PA-B2G1(.00) 0.004s 0.005s 0.015s 0.016s 0.061s 0.084s
PA-B2G1(.04) 0.018s 0.022s 0.054s 0.089s 0.193s 0.306s
PA-B2G1(.08) 0.043s 0.049s 0.148s 0.222s 0.520s 0.871s
PA-B2G2(.00) 0.005s 0.007s 0.025s 0.029s 0.084s 0.119s
PA-B2G2(.04) 0.020s 0.025s 0.071s 0.074s 0.217s 0.264s
PA-B2G2(.08) 0.046s 0.057s 0.171s 0.177s 0.725s 0.784s
@@
Table 5: Average time required to generate noise $\mathbf{g}_{s}$ of different sizes using PA-B2G in modes I and $I$ with varying $\Delta_{g}$.
5.5 Computational Complexity
@@
This section evaluates the computational complexity of PA-B2G in modes I and II. The complete message-hiding pipeline encompasses four stages: mapping the secret bit sequence $\mathbf{b}$ to noise $\mathbf{g}_{s}$, reverse ODE solving, forward ODE solving, and recovering $\mathbf{b}$ from $\mathbf{g}_{s}$. Table˜5 presents the average execution time for the initial mapping stage, with the error tolerance $e$ set to $0.0185$. While the computational overhead scales with noise dimensionality and $\Delta_{g}$, it can be mitigated by increasing the correction step $\Delta_{c}$ or error tolerance $e$. Notably, PA-B2G consistently completes this initial mapping in one second, regardless of the mode or noise dimensions. While the time consumption of the ODE solving processes also depends on specific diffusion architectures and sampling steps, it is considered beyond the scope of this paper. The time required for message recovery is marginal compared to the mapping process and is thus omitted.
6 Conclusion

### 04 Mar 2026 v5

@@ 5.3. Method Comparison
[…]
Comparison with Prior DM-GIS Methods. We compare PA-B2G with message mapping methods including MN, MB, MC, and Multi-bits proposed in (Kim et al., 2025), GSD (Wei et al., 2023), and Gaussian Shading (GS) (Yang et al., 2024). All message mapping methods are evaluated on CIFAR-10 $32\times 32$, FFHQ $64\times 64$, and LSUN-Bedroom $256\times 256$ datasets. For each dataset, we equip each mapping method with the same generative backbone and sampling scheduler. Specifically, we employ the $2^{\text{nd}}$ order Heun ODE solver with 20 sampling steps on CIFAR-10 $32\times 32$ and with 42 sampling steps on FFHQ $64\times 64$, while using the $1^{\text{st}}$ order ODE solver proposed in DDIM (Song et al., 2020a) with 18 sampling steps on LSUN-Bedroom $256\times 256$. We set the error tolerance $e=0.0185$ defined in Algorithm 3 for CIFAR-10 $32\times 32$ and FFHQ $64\times 64$, and $e=0.007$ for LSUN-Bedroom $256\times 256$. Before message hiding, we pad all secret bits and then encrypt them to meet required payloads. All mapping methods utilize the same random seed for stego image generation, and the corresponding generated images are saved in PNG format. Note that the payloads listed in Section 5.3 are effective payloads.
@@
Figure 3. Visualization results of the adjustment strategies adopted by (Kim et al., 2025) and PA-B2G on CIFAR-10, FFHQ, and LSUN-Bedroom.
[…]
Comparison with GAN- and Flow-based Methods. We also compare PA-B2G with two flow-based methods, GSF (Wei et al., 2022c) and S2IRT (Zhou et al., 2022), and a GAN-based method, GSN (Wei et al., 2022b), on the CelebA $64\times 64$ dataset with payloads ranging from 3 to 9 bpp. GSF and S2IRT utilize the same pre-trained Glow model  [^1] https://github.com/rosinality/glow-pytorch. For S2IRT, to maintain computational efficiency, we set the group number $K$ to 30 for a 3 bpp payload and to 40 for 6 bpp. At higher payloads, S2IRT requires a larger $K$, making it too time-consuming to generate 50,000 stego images; thus, we do not report its corresponding results in Section 5.3. Given that PA-B2G in Mode I is superior to Mode II in extraction accuracy, we select it with $\Delta_{g}=0.08$ and $e=0.007$ for comparison. We employ the pre-trained diffusion model from DDIM (Song et al., 2020a) but replace its sampling scheduler with the Heun ODE solver, as suggested in Section 4.4. As shown in Section 5.3, PA-B2G significantly outperforms GAN- and flow-based methods, particularly in image quality and message extraction accuracy, primarily due to the powerful generation capabilities of diffusion models and our provable, adjustable message mapping. Regarding security, S2IRT achieves the best anti-detectability, as the extra perturbation key it introduces ensures that the generated noise strictly follows a pure Gaussian distribution. In contrast, GSF and GSN alter the noise distribution and are thus easily detected by UCNet.
@@
Figure 4. Visualization results of combining PA-B2G in mode I with Stable Diffusion on text-to-image and image-to-image tasks.
@@
Comparison with Prior Adjustment Strategies. Kim et al. (Kim et al., 2025) identify a trade-off between stego image quality and extraction accuracy, proposing MN, MB, and MC as balancing options. To evaluate our adjustment strategy (Algorithm 3), Section 5.3 reports the FID score, $\overline{Acc}$, and $Acc_{s}$ for these methods alongside PA-B2G (modes I &\& II) across varying $\Delta_{g}$. For better intuition, Figure 3 illustrates further adjustment results, confirming that image quality, security, and accuracy can be effectively balanced by tuning $\Delta_{g}$. This supports our analysis in Section 4.1 and reveals that Mode I provides a more natural, fine-grained adjustment than (Kim et al., 2025). By modulating $\Delta_{g}$, PA-B2G can readily match the performance of MN, MB, and MC, as they represent special cases of our generalized framework.
5.4. Diffusion Model Watermarking
[…]
@@ 5.4. Diffusion Model Watermarking
Robustness evaluation of average message extraction accuracy (%) under various image distortions (JPEG compression with different quality factors, random cropping with varying ratios, Gaussian noise with varying standard deviations, and Gaussian blurring with varying radii) across different payload sizes (256, 512, and 1024 bits) JPEG Random Crop Gaussian Noise Gaussian Blur Payload 90 75 50 0.1 0.3 0.5 0.01 0.03 0.05 2 3 4 256 bits 100.00% 99.89% 99.46% 99.98% 99.04% 87.05% 99.85% 98.03% 94.45% 99.94% 99.25% 96.64% 512 bits 99.80% 99.27% 98.20% 99.51% 94.20% 78.25% 98.79% 94.95% 85.58% 99.26% 96.58% 91.48% 1024 bits 98.60% 97.30% 94.45% 97.48% 87.88% 71.22% 96.64% 88.86% 81.93% 96.04% 90.76% 83.43%
@@
In this section, we further validate the effectiveness of PA-B2G in model I for watermarking diffusion models. We employ Dreamshaper 7, which is fine-tuned on Stable Diffusion 1.5, as the generative backbone, and utilize a first-order DDIM solver with 50 sampling steps for stego image generation. We evaluate three effective payloads—256, 512, and 1024 bits—and report their corresponding average message extraction accuracy ($\overline{Acc}$) over 1000 stego images against various lossy processing techniques in Section 5.4. These techniques include JPEG compression with different quality factors, random cropping with varying ratios, Gaussian noise with varying standard deviations, and Gaussian blurring with varying radii. For PA-B2G, the error tolerance $e$ is set to 0.02, code length $l$ to 1, and $\Delta_{g}$ to 0.05.
@@
As reported in Section 5.4, PA-B2G exhibits robust resistance to nearly all lossy processing when the effective payload is 256 bits. Notably, even with a cropping ratio of 0.5, PA-B2G still guarantees an extraction accuracy of 87.05%. While the extraction accuracy gradually decreases as the payload increases, PA-B2G with a 1024-bit payload still effectively resists JPEG compression ($\overline{Acc}>94\%$), which is a common operation on social media platforms. In addition to extraction accuracy, we present the visual generation results for text-to-image and image-to-image tasks in Figure 4. As shown, the stego images produced by PA-B2G exhibit visual quality indistinguishable from normally generated images. These experimental results demonstrate that PA-B2G can be effectively extended to the watermarking of diffusion models.
[…]
@@
Average time required to generate noise $\mathbf{g}_{s}$ of different sizes using PA-B2G in modes I and II with varying $\Delta_{g}$. $32\!\times\!32\!\times 3$ $64\!\times\!64\!\times 3$ $128\!\times\!128\!\times 3$ Method 3 bpp 6 bpp 3 bpp 6 bpp 3 bpp 6 bpp PA-B2G1(.00) 0.004s 0.005s 0.015s 0.016s 0.061s 0.084s PA-B2G1(.04) 0.018s 0.022s 0.054s 0.089s 0.193s 0.306s PA-B2G1(.08) 0.043s 0.049s 0.148s 0.222s 0.520s 0.871s PA-B2G2(.00) 0.005s 0.007s 0.025s 0.029s 0.084s 0.119s PA-B2G2(.04) 0.020s 0.025s 0.071s 0.074s 0.217s 0.264s PA-B2G2(.08) 0.046s 0.057s 0.171s 0.177s 0.725s 0.784s
5.5. Computational Complexity
@@
This section evaluates the computational complexity of PA-B2G in modes I and II. The complete message-hiding pipeline encompasses four stages: mapping the secret bit sequence $\mathbf{b}$ to noise $\mathbf{g}_{s}$, reverse ODE solving, forward ODE solving, and recovering $\mathbf{b}$ from $\mathbf{g}_{s}$. Section 5.4 presents the average execution time for the initial mapping stage, with the error tolerance $e$ set to $0.0185$. While the computational overhead scales with noise dimensionality and $\Delta_{g}$, it can be mitigated by increasing the correction step $\Delta_{c}$ or error tolerance $e$. Notably, PA-B2G consistently completes this initial mapping in one second, regardless of the mode or noise dimensions. While the time consumption of the ODE solving processes also depends on specific diffusion architectures and sampling steps, it is considered beyond the scope of this paper. The time required for message recovery is marginal compared to the mapping process and is thus omitted.
6. Conclusion